# Reconstruction Alignment Improves Unified Multimodal Models

### v2

@@ Abstract
Unified multimodal models (UMMs) unify visual understanding and generation within a single architecture. However, conventional training relies on image–text pairs (or sequences) whose captions are typically sparse and miss fine-grained visual details—even when they use hundreds of words to describe a simple image. We introduce Reconstruction Alignment (RecA), a resource-efficient post-training method that leverages visual understanding encoder embeddings as dense “text prompts,” providing rich supervision without captions. Concretely, RecA conditions a UMM on its own visual understanding embeddings and optimizes it to reconstruct the input image with a self-supervised reconstruction loss, thereby realigning understanding and generation. Despite its simplicity, RecA is broadly applicable: across autoregressive, masked-autoregressive, and diffusion-based UMMs, it consistently improves generation and editing fidelity. With only 27 GPU-hours, post-training with RecA substantially improves image generation performance on GenEval (0.73→0.90) and DPGBench (80.93→88.15), while also boosting editing benchmarks (ImgEdit 3.38→3.75, GEdit 6.94→7.27). Notably, RecA surpasses much larger open-source models and applies broadly across diverse UMM architectures, establishing it as an efficient and general post-training alignment strategy for UMMs.
Figure 1: Post-training UMMs with reconstruction alignment (i.e., RecA) substantially improves image generation and editing. Left: performance comparison on GenEval and DPGBench, where a 1.5B-parameter model post-trained with RecA surpasses much larger models across multiple benchmarks (Table 1: GenEval, DPGBench and Wise; Table 3: ImgEdit and GEdit-Bench-EN); Middle: compared with GPT-4o, RecA follows generation instructions more faithfully, especially for color attributes and spatial positions; Right: for editing, RecA better preserves instance identity, overall layout, and object shapes of the original images, such as the girl’s lips.
@@ 1 Introduction
Figure 2: Dense supervision from visual embeddings. a) Typical image generation models are trained on image–caption pairs and/or sequences whose text is a sparse representation of visual information. An image is worth far more than a hundred of words and contains rich details that text alone cannot capture. As shown in the left three examples, even lengthy captions (500 words) miss key aspects such as textures, styles, layouts, shapes, and attributes, leading to imperfect generations relative to the original image. b) By contrast, embeddings from visual understanding encoders, e.g., CLIP, preserve richer and more faithful semantics. Can these image–embedding pairs provide the dense supervision needed to enhance image generation and editing? Surprisingly, the answer is yes: we find that image–embedding pairs can improve T2I and image editing in a zero-shot manner.
Building on the success of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Yang et al., 2024a), researchers have developed Multimodal Large Language Models (MLLMs) (Liu et al., 2024b; Bai et al., 2023; 2025; Radford et al., 2021; Zhai et al., 2023; Chen et al., 2024; Zhu et al., 2025) with strong visual understanding performance. Recently, unified multimodal models (UMMs), or Omni Models, have been proposed to both understand and generate across modalities—reading and writing visual content and text within a single architecture (Tong et al., 2024a; Ge et al., 2024; Wu et al., 2024b; Chen et al., 2025c; Pan et al., 2025a; OpenAI, 2024; Li et al., 2025; Wu et al., 2025a). The academic community envisions that a unified framework can inherit the reasoning and world knowledge of LLMs while extending them to content generation.
@@
However, UMMs face a fundamental limitation: conventional training relies on image–text pairs, where captions provide supervision. Even captions spanning hundreds of words omit critical visual details such as spatial layout, geometry, or fine-grained attributes (Figure 2), introducing systematic biases into the learned representations. For instance, since captions rarely describe broccoli’s color, models tend to overfit to the rule broccoli $\to$ green, often collapsing to green outputs or failing on atypical prompts like “a yellow broccoli” (Figure 3). This misalignment motivates us to explore alternative forms of supervision. Rather than relying on captions, we leverage embeddings from visual understanding encoders (Radford et al., 2021; Zhai et al., 2023; Chen et al., 2024; Zhu et al., 2025), which map pixels into a language-aligned semantic space interpretable by unified multimodal models. Crucially, embeddings from understanding encoders (e.g., CLIP, SigLIP) capture semantic structure far more effectively than those from generation encoders (e.g., VAE, VQ-GAN). These semantic embeddings provide dense, semantically grounded supervision without paired captions, raising a central question: Can we improve the generation capabilities of UMMs by training them with semantic embeddings as maximally informative “text prompts”?
Building on these insights, we propose RecA, a resource-efficient post-training strategy. The core idea is simple: condition UMMs on their own understanding vision encoder embeddings—dense “visual prompts” that encode layout, color, and attributes beyond what captions capture—and train them to reconstruct the image. This semantic reconstruction provides richer supervision without additional labels. Despite its simplicity, RecA yields substantial improvements. With only 27 A100 GPU hours, a 1.5B-parameter UMM post-trained with RecA surpasses GPT-4o and significantly larger open-source models, achieving state-of-the-art results on GenEval (0.86) and DPGBench (87.21). Importantly, these gains are obtained without any GPT-4o-Image distillation data or reinforcement learning (OpenAI, 2024; Chen et al., 2025a; b), in contrast to prior work. Moreover, when post-trained with GPT-4o data, RecA further improves to GenEval (0.90) and DPGBench (88.15), substantially surpassing previous methods that use much larger model sizes. RecA also boosts image editing quality, raising ImgEdit from 3.38→3.75 and GEdit from 6.94→7.27. Moreover, RecA applies broadly across UMM families—including Show-o (Xie et al., 2025b) (AR), Harmon (Wu et al., 2025d) (AR+MAR), OpenUni (Wu et al., 2025c) (AR+Diffusion), and BAGEL (Deng et al., 2025) (AR+Diffusion)—highlighting its generality.
@@
  • Generality: We show that RecA consistently improves diverse UMM architectures, from autoregressive to hybrid frameworks.
  • Performance: We demonstrate strong empirical gains, with a 1.5B-parameter model surpassing GPT-4o and larger open-source models using only 27 A100 GPU hours, significantly outperforming prior state of the art without distillation or reinforcement learning (RL).
Figure 4: Overview of the semantic reconstruction alignment (RecA) pipeline. A visual understanding encoder (e.g., CLIP or DINO) extracts semantic features from the input image, which are fused with template text embeddings and passed to a Unified Multimodal Model (UMM) to regenerate the image. The UMM is optimized with a self-supervised loss (diffusion or cross-entropy) between the original and reconstructed images or image latents. We freeze the visual understanding encoder, except in cases where the UMM employs the same encoder for both understanding and generation (e.g., Harmon, which uses an MAE encoder for both tasks). At inference time, RecA requires no additional inputs, operating as a standard UMM.
@@ 2 Reconstruction Alignment
In this section, we present Reconstruction Alignment (RecA) as a self-supervised image reconstruction objective. By training the model to reconstruct images from its visual understanding encoder embeddings, RecA provides dense supervision that captures fine-grained visual details often omitted by text captions. We show the overall pipeline at Figure 4.
Figure 5: Post-training with RecA restores visual details beyond the baseline. For each query image (left), we feed its visual understanding embeddings back into the UMM with the instruction “Describe the image in detail.” The baseline model (centre)’s visual responses, i.e., images, preserve the main subject but distort layout, textures, and colors, while RecA markedly restores visual details like geometry, color, and overall fidelity.
@@ 2.1 Motivation and Setup
Since the visual understanding encoder projects image features into a language-aligned semantic space, a truly "unified" model of visual understanding and generation should be capable of leveraging dense, rich semantic information and generating the same image. To test this, we extract semantic embeddings from the visual understanding encoder, insert them into a prompt template (e.g., “Describe the image in detail.”), and ask the UMM to regenerate the input image.
As shown in Figure 5, the results are revealing: current UMMs preserve the main subject but scramble spatial layout and composition. This suggests that UMMs trained under conventional paradigms fail to fully exploit the model’s fine-grained semantic, i.e., the latent spaces for understanding and generation remain only partially aligned.
@@ 2.2 RecA Training Paradigm
where $\mathcal{L}(\cdot,\cdot)$ denotes the training loss (e.g., cross-entropy for autoregressive models (Xie et al., 2025b; Chen et al., 2025d), diffusion loss for diffusion-based models (Zhou et al., 2025; Deng et al., 2025)), $t_{\text{prompt}},t_{\text{question}},t_{\text{answer}}$ are text inputs/outputs, $\mathbf{h}_{v}$ are embeddings extracted from the visual understanding encoder, and $I_{\text{gt}}$ is the ground-truth image (we omit the explicit VAE decoder notation for simplicity). $f_{\theta}(\cdot)$ represents the UMM parameters $\theta$.
Our key idea is to replace conventional T2I supervision with an image reconstruction loss. Instead of using text captions that are inherently sparse in visual information, we condition the UMM on its own visual understanding embeddings, which provide richer semantics. The reconstruction loss is:
@@
where $t_{\text{template}}$ is a simple prompt template triggering image reconstruction. The overall training objective can be formulated as:
@@
Training configuration. We set $\lambda_{\text{{RecA}}}=1,\lambda_{\text{t2i}}=0$. For UMMs that share parameters between understanding and generation, we set $\lambda_{\text{i2t}}=1$ to preserve image-to-text capability; otherwise, we set $\lambda_{\text{i2t}}=0$ and freeze the understanding component. We freeze the visual understanding encoder in most cases, except when the UMM employs the same encoder for both tasks. Input images are resized to the minimum resolution accepted by the visual encoder.
Model inference. At inference time, our post-trained UMM operates identically to a standard UMM and requires no additional visual embeddings. For image generation, only a text prompt is needed; for image editing, the inputs remain the text prompt and the original image. This preserves the model’s original usability while providing enhanced generation capabilities.
@@ 2.3 Difference Between RecA and Previous Works
RecA introduces semantic-level image reconstruction as a native post-training objective for UMMs, directly leveraging visual understanding priors to improve both generation and editing, without auxiliary modules or text-image supervision. In contrast, previous works integrate reconstruction in different ways: (I) diffusion-supervised enhancement (e.g., DEEM, DIVA, ViLex) (Luo et al., 2024; Wang et al., 2024b; 2025d; Ma et al., 2025a) leverages pretrained diffusion models to regularize vision encoders and improve understanding; (II) reconstruction from hidden states (e.g., ROSS) (Wang et al., 2024a; 2025b)) adds lightweight decoders to regenerate input images from intermediate embeddings, thereby regularizes the model to preserve fine-grained details for VLMs; (III) representation alignment (e.g., REPA) (Yu et al., 2024) introduces additional alignment modules that map noisy hidden states from the VAE in denoising networks to clean image representations obtained from external, pretrained visual encoders; and (IV) reconstruction as a prior (e.g., Lumos) (Ma et al., 2025b) adds additional DINO features (Caron et al., 2021) into the attention blocks of the diffusion model, which is further trained on large-scale text-to-image data. Our method is fundamentally different in terms of the methodology, architecture, motivation and task.
2.4 Difference Between RecA and Classifier-Free Guidance (CFG)
Classifier-free guidance (CFG (Ho & Salimans, 2022)) is typically used to improve fidelity. At each generation step, we compute a conditional prediction $\mathbf{o}_{\text{cond}}$ and an unconditional prediction $\mathbf{o}_{\text{uncond}}$, where $\mathbf{o}$ denotes either the autoregressive head’s logits (Chen et al., 2025d; Xie et al., 2025b) or the diffusion head’s predicted noise (Zhou et al., 2025; Deng et al., 2025). The output is given by:

### v3

@@ Abstract
Unified multimodal models (UMMs) unify visual understanding and generation within a single architecture. However, conventional training relies on image–text pairs (or sequences) whose captions are typically sparse and miss fine-grained visual details, even when they use hundreds of words to describe a simple image. We introduce Reconstruction Alignment (RecA), a resource-efficient post-training method that leverages visual understanding encoder embeddings as dense “text prompts,” providing rich supervision without captions. Concretely, RecA conditions a UMM on its own visual understanding embeddings and optimizes it to reconstruct the input image with a self-supervised reconstruction loss, thereby realigning understanding and generation. Despite its simplicity, RecA is broadly applicable: across autoregressive, masked-autoregressive, and diffusion-based UMMs, it consistently improves generation and editing fidelity. With only 27 GPU hours, post-training with RecA substantially improves image generation performance on GenEval (0.73→0.90) and DPGBench (80.93→88.15), while also boosting editing benchmarks (ImgEdit 3.38→3.75, GEdit 6.94→7.27). Notably, RecA surpasses much larger open-source models and applies broadly across diverse UMM architectures, establishing it as an efficient and general post-training alignment strategy for UMMs.
Figure 1: Post-training UMMs with reconstruction alignment (i.e., RecA) substantially improve image generation and editing. Left: performance comparison on GenEval and DPGBench, where a 1.5B-parameter model post-trained with RecA surpasses much larger models across multiple benchmarks (Table 1: GenEval, DPGBench and WISE; Middle: compared with GPT-4o, RecA follows generation instructions more faithfully, especially for color attributes and spatial positions; Right: for editing, RecA better preserves instance identity, overall layout, and object shapes of the original images, such as the girl’s lips.
@@ 1 Introduction
Figure 2: Dense supervision from visual embeddings. a) Typical image generation models are trained on image–caption pairs and/or sequences whose text is a sparse representation of visual information. An image is worth far more than a hundred words and contains rich details that text alone cannot capture. As shown in the left three examples, even lengthy captions (500 words) miss key aspects such as textures, styles, layouts, shapes, and attributes, leading to imperfect generations relative to the original image. b) By contrast, embeddings from visual understanding encoders, e.g., CLIP, preserve richer and more faithful semantics. Can these image–embedding pairs provide the dense supervision needed to enhance image generation and editing? Surprisingly, the answer is yes: we find that image–embedding pairs can improve T2I and image editing in a zero-shot manner.
Building on the success of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Yang et al., 2024a), researchers have developed Multimodal Large Language Models (MLLMs) (Liu et al., 2024b; Bai et al., 2023; 2025; Radford et al., 2021; Zhai et al., 2023; Chen et al., 2024b; Zhu et al., 2025) with strong visual understanding performance. Recently, unified multimodal models (UMMs), or Omni Models, have been proposed to both understand and generate across modalities—reading and writing visual content and text within a single architecture (Tong et al., 2024a; Shi et al., 2024; Zhang et al., 2025a; Ge et al., 2024; Wu et al., 2024b; Pan et al., 2025a; OpenAI, 2024; Li et al., 2025; Wu et al., 2025a). The academic community envisions that a unified framework can inherit the reasoning and world knowledge of LLMs while extending them to image generation.
@@
However, UMMs face a fundamental limitation: conventional training relies on image–text pairs, where text captions provide supervision. Even captions spanning hundreds of words omit critical visual details such as spatial layout, geometry, or fine-grained attributes (Figure 2), introducing systematic biases into the learned representations. For instance, because captions rarely describe broccoli’s color, models tend to overfit to the rule broccoli $\to$ green, often collapsing to green outputs when given atypical prompts like “a yellow broccoli” (Figure 3). This misalignment motivates us to explore alternative forms of supervision. Rather than relying on captions, we leverage embeddings from UMM’s visual understanding encoders (Radford et al., 2021; Zhai et al., 2023; Chen et al., 2024b; Zhu et al., 2025), which can map pixels into a language-aligned semantic space interpretable by themselves. Crucially, embeddings from understanding encoders (e.g., CLIP, SigLIP) capture semantic structure far more effectively than those from generation encoders (e.g., VAE, VQ-GAN). These semantic embeddings provide dense, semantically grounded supervision without paired captions, raising a central question: Can we improve the generation capabilities of UMMs by training them with semantic embeddings as maximally informative “text prompts”?
Building on these insights, we propose RecA, a resource-efficient post-training strategy. The core idea is simple: condition UMMs on their own visual understanding encoder embeddings—dense “visual prompts” that encode layout, color, and attributes beyond what captions capture—and train them to reconstruct the image. This semantic reconstruction provides richer supervision without additional labels. Despite its simplicity, RecA yields substantial improvements. With only 27 A100 GPU hours, a 1.5B-parameter UMM post-trained with RecA surpasses significantly larger open-source models, achieving state-of-the-art results on GenEval (0.86) and DPGBench (87.21). Importantly, these gains are obtained without any GPT-4o-Image distillation data or reinforcement learning (OpenAI, 2024; Chen et al., 2025a; b), in contrast to prior work. Moreover, when post-trained with GPT-4o-Image data, RecA further improves to GenEval (0.90) and DPGBench (88.15). RecA also boosts image editing quality, raising ImgEdit from 3.38 to 3.75 and GEdit from 6.94 to 7.27. Moreover, RecA applies broadly across UMM families, including Show-o (Xie et al., 2025b) (AR), Harmon (Wu et al., 2025d) (AR+MAR), OpenUni (Wu et al., 2025c) (AR+Diffusion), and BAGEL (Deng et al., 2025) (AR+Diffusion), highlighting its generality.
@@
  • Generality: We show that RecA consistently improves diverse UMM architectures, spanning autoregressive models to hybrid frameworks that integrate autoregressive and diffusion models.
  • Performance: We demonstrate strong empirical gains, with a 1.5B-parameter model surpassing GPT-4o and larger open-source models using only 27 A100 GPU hours, significantly outperforming prior state-of-the-art methods without distillation or reinforcement learning (RL).
Figure 4: Overview of the semantic reconstruction alignment (RecA) pipeline. A visual understanding encoder (e.g., CLIP or SigLIP) extracts semantic features from the input image, which are fused with template text embeddings and passed to the Unified Multimodal Model (UMM) to regenerate the image. The UMM is optimized with a self-supervised loss (diffusion or cross-entropy) between the original and generated images or image latents. We freeze the understanding encoder except in cases where the UMM employs shared encoder for both understanding and generation (e.g., Harmon). At inference time, RecA requires no additional inputs, operating as a standard UMM.
@@ 2 Reconstruction Alignment
In this section, we present Reconstruction Alignment (RecA) as a self-supervised image reconstruction objective. By training the model to reconstruct images from its visual understanding embeddings, RecA provides dense supervision that captures fine-grained visual details often omitted by text captions. We show the overall pipeline in Figure 4.
Figure 5: Post-training with RecA restores visual details missed by the baseline models. For each query image (left), we feed its visual understanding embeddings back into the UMM with the instruction “Describe the image in detail.” The baseline model (center)’s visual responses, i.e., images, preserve the main subject but distort layout, textures, and colors, while RecA markedly restores visual details like geometry, color, and overall fidelity.
@@ 2.1 Motivation and Setup
Equipped with a visual understanding encoder, modern multimodal LLMs (MLLMs) effectively treat images as rich, dense context and successfully handle not only basic perceptual and question answering tasks, but also challenging multimodal reasoning problems such as symbolic, spatial, and mathematical reasoning (Bai et al., 2025; OpenAI, 2024; Huang et al., 2025; Chen et al., 2024a).
One may expect that this dense, text-like context can be naturally extended to the image generation task. However, when we extract embeddings from the understanding encoder, insert them into a prompt template (e.g., “Describe the image in detail.”), and ask the UMM to generate, the results in Figure 5 reveal a gap: the main subject is preserved, but spatial layout and composition are scrambled. This raises the hypothesis that the image itself can be treated as a dense “text prompt” to train the model, enabling the generation branch of the UMM to better exploit the LLM features and improve its generation capability.
@@ 2.2 RecA Training Paradigm
where $\mathcal{L}(\cdot,\cdot)$ denotes the training loss (e.g., cross-entropy for autoregressive models (Xie et al., 2025b; Chen et al., 2025d), diffusion loss for diffusion-based models (Zhou et al., 2025; Deng et al., 2025)), $t_{\text{prompt}},t_{\text{question}},t_{\text{answer}}$ are text inputs/outputs, $\mathbf{h}_{v}$ are embeddings extracted from the understanding encoder, and $I_{\text{gt}}$ is the ground-truth image (we omit the explicit VAE decoder notation for simplicity). $f_{\theta}(\cdot)$ represents the UMM with parameters $\theta$. Preliminaries of different image generation paradigms are listed in Appendix A.
Our key idea is to replace conventional T2I supervision with a semantic-level image reconstruction loss. Instead of using text captions that are inherently sparse in visual information, we condition the UMM on its own visual understanding embeddings. The reconstruction loss is:
@@
where $t_{\text{template}}$ is a prompt template triggering image reconstruction (e.g., Describe the image in detail). More details of the templates are provided in Appendix C. The overall training loss is:
@@
We set $\lambda_{\text{{RecA}}}=1,\lambda_{\text{t2i}}=0$. For UMMs that share parameters between understanding and generation, we set $\lambda_{\text{i2t}}=1$ to preserve I2T capability; otherwise, we set $\lambda_{\text{i2t}}=0$ and freeze the understanding component. We freeze the understanding encoder in most cases, except when the UMM employs the same encoder for both tasks (e.g., Harmon (Wu et al., 2025d)).
Input resolution. Previous works show that higher-resolution visual understanding embeddings retain substantially more pixel-level detail (Allakhverdov et al., 2025; Lin et al., 2025). To encourage the model to focus on semantic-level reconstruction, we resize input images to the minimum resolution accepted by the understanding encoder.
Model inference. At inference time, our post-trained UMM operates identically to a standard one and requires no additional visual embeddings. For image generation, only a text prompt is needed; for image editing, the inputs remain the text prompt and the original image. This preserves the model’s original usability while providing enhanced generation capabilities.
@@ 2.3 Difference Between RecA and Previous Works
Previous works integrate image reconstruction in different ways: (I) diffusion-supervised enhancement (e.g., ViLex) (Wang et al., 2025d; Ma et al., 2025a; Luo et al., 2024; Wang et al., 2024b) leverages pretrained diffusion models to regularize vision encoders and improve visual understanding; (II) reconstruction from hidden states (e.g., ROSS) (Wang et al., 2024a; 2025b)) adds lightweight decoders to reconstruct input images from intermediate embeddings, thereby regularizing the model to preserve fine-grained details for visual understanding; (III) representation alignment (e.g., REPA, VA-VAE) (Yu et al., 2024; Yao et al., 2025) introduces additional alignment modules that map hidden states in DiT (Peebles & Xie, 2022; Ma et al., 2024a) or VAE (Kingma & Welling, 2013) to representations obtained from external, pretrained vision encoders; and (IV) reconstruction as a prior (e.g., Lumos) (Ma et al., 2025b) adds additional DINO features (Caron et al., 2021) into the attention blocks of the diffusion model, which is further trained on large-scale text-to-image data.
RecA treats images as dense prompts and adopts semantic-level reconstruction as a native post-training objective for UMMs, requiring no auxiliary modules or text-to-image data. RecA is fundamentally different in terms of the methodology, architecture, motivation and task.
@@ 3 Experiments
Table 1: Results on GenEval and DPGBench. Scores marked with (*) are our reproduced results using 12 random seeds. We use Harmon-1.5B as baseline and post-train it with RecA. The gray-colored rows denote private models, and their results are cited from (Yan et al., 2025; Geng et al., 2025). Arrows ($\uparrow$) denote that higher is better. Detailed comparison is listed in Appendix E.3.
[…]
Harmon* 1.5B 0.99 0.87 0.69 0.86 0.45 0.51 0.73 80.93
SD3-Medium 2B 0.99 0.94 0.72 0.89 0.33 0.60 0.74 84.08
Janus-Pro 7B 0.99 0.89 0.59 0.90 0.79 0.66 0.80 84.33
FLUX-dev 12B 0.99 0.85 0.74 0.79 0.21 0.48 0.68 84.00
BAGEL* 14B 0.99 0.93 0.80 0.86 0.51 0.63 0.79 84.03
Playground-v3 24B 0.99 0.95 0.72 0.82 0.50 0.54 0.76 87.06
GPT-4o-Image - 0.99 0.92 0.85 0.89 0.74 0.71 0.84 86.23
RecA 1.5B 1.00 0.98 0.71 0.93 0.76 0.77 0.86 87.21
Models Trained with GPT-4o Data
Ovis-U1 3.6B 0.98 0.98 0.90 0.92 0.79 0.75 0.89 83.72
OmniGen2 7B 1.00 0.95 0.64 0.88 0.55 0.76 0.80 83.57
BLIP3-o* 8B 1.00 0.92 0.63 0.91 0.86 0.67 0.83 80.73
RecA 1.5B 1.00 0.97 0.76 0.94 0.91 0.83 0.90 88.15